# F2PAD: A General Optimization Framework for Feature-Level to Pixel-Level Anomaly Detection

### v1

@@ Abstract
Image-based inspection systems have been widely deployed in manufacturing production lines. Due to the scarcity of defective samples, unsupervised anomaly detection that only leverages normal samples during training to detect various defects is popular. Existing feature-based methods, utilizing deep features from pretrained neural networks, show their impressive performance in anomaly localization and the low demand for the sample size for training. However, the detected anomalous regions of these methods always exhibit inaccurate boundaries, which impedes the downstream tasks. This deficiency is caused: (i) The decreased resolution of high-level features compared with the original image, and (ii) The mixture of adjacent normal and anomalous pixels during feature extraction. To address them, we propose a novel unified optimization framework (F2PAD) that leverages the Feature-level information to guide the optimization process for Pixel-level Anomaly Detection in the inference stage. The proposed framework is universal and can enhance various feature-based methods with limited assumptions. Case studies are provided to demonstrate the effectiveness of our strategy, particularly when applied to three popular backbone methods: PaDiM, CFLOW-AD, and PatchCore. Our source code and data will be available upon acceptance.
Note to Practitioners
Visual quality inspection systems are popular in modern industries for defect detection on the surface of manufacturing products. Since the defects are rare during production, it is suitable to collect the normal samples as the training dataset and then detect unseen defects at testing. The feature-based methods are popular, which can localize the anomaly well and handle small-batch production, i.e., only a few normal samples are collected for training. However, these methods cannot guarantee precise anomaly boundaries, which is undesired for anomaly characterization and further diagnosis. This paper analyzes the two issues that correspond to the above problem and proposes a novel general optimization framework to achieve precise anomaly boundary segmentation at the pixel level. The proposed method can be applied to enhance a wide range of existing feature-based methods, including the well-known PaDiM, CFlow, and PatchCore, which is demonstrated by the extensive case studies. By implementing our method, practitioners can obtain more accurate anomaly information, benefiting other downstream tasks and objectives.
@@
Image anomaly detection, unsupervised learning, anomaly segmentation, optimization, feature-based anomaly detection, neural networks.
@@ I Introduction
Automated visual inspection of products by deploying cost-efficient image sensors has become a critical concern in industrial applications in recent years, ranging from microscopic fractions and crack detection on electrical commutators [1] to fabric defect detection on textiles [2]. In realistic production lines, defective products are always scarce. Moreover, all possible types of defects that occurred during production are impossible to know beforehand [3, 4]. In this scenario, only utilizing normal (non-defective) samples to train modern machine learning models for defect detection becomes essential. Therefore, unsupervised anomaly detection has become a mainstream setting in practical applications.
Besides determining whether a sample is anomalous or not, it is also essential to precisely localize and segment the defective regions due to the following reasons: (i) It enables the characterization of surface defects, such as shape, size, and orientation [5], which contributes to the evaluation of product quality [6]. (ii) It benefits the subsequent defect classification and recognition [7], as the models could focus more on the discriminative regions. An example is provided in Fig. 1, where the target defective region is masked as white in the ground truth (GT).
@@ II-A Image Reconstruction-Based Methods
Reconstruction-based image anomaly detection methods aim to reconstruct the normal samples accurately during training, assuming imperfect reconstruction of defective regions for an anomalous image during inference. The pixel differences between the input and reconstructed images can be directly transformed into anomaly scores and further thresholded as a binary anomaly mask. Therefore, the key research goal is to pursue a more accurate reconstruction. Autoencoder (AE) is the most prevalent architecture for image reconstruction [23, 24, 25]. In AEs, a common choice of the backbone network is CNN. Vision transformers were introduced in [26, 27] to capture more global structural information. Besides the AEs, GANs [11, 28] and diffusion models [29, 30] are also popular choices.
Instead of focusing on the architecture of the reconstruction network, there are various strategies to improve the reconstruction accuracy. For example, the structure similarity index measure (SSIM) was introduced in [9] to incorporate the perceptual difference of anomaly. Additionally, to address the deterioration of image reconstruction caused by anomalies, [31, 32] enforced the reconstructed image on the learned manifold. Alternatively, [33] adopted a self-supervised learning way by generating synthetic anomalies to avoid reconstructing undesired anomalies.
@@ II-B Feature-Based Methods
PaDiM [13] adopted the multivariate Gaussian distributions to model the normal features. To allow more complicated probabilistic distributions, DifferNet [15] adopted a neural network, i.e., normalizing flow, to estimate the density of features. Furthermore, CFLOW-AD [16] improved the DifferNet by modeling the conditional probabilities of features on their locations. Similarly, a 2D normalizing flow was proposed in [34] to retain the spatial relationship between features. The memory bank-based method that collects all normal features was first proposed in [35], while the average distance of a test feature to its k-nearest neighbors was regarded as the anomaly score. To improve the inference speed, the coreset technique was adopted in PatchCore[19] by greedily selecting the farthest features to construct a reduced memory bank. Similarly, [36] proposed a gradient preference strategy to determine the most significant features for anomaly detection. In addition, the deep support vector domain description (SVDD) was adopted in [22, 37] to map normal image patches within compact spheres. Teacher-student networks were also popular in literature [38], which effectively transfer normal features from the teacher network to the student model but fail for anomalous features.
One issue of the above methods is the decreased resolution of the feature map, as stated in Issue 1. To tackle it, the work [17] learned dense features in the image size by using the multipooling and unwarping strategy in [39]. However, it is only applicable to local patch descriptions by CNNs. Furthermore, the anomaly scores were defined as the gradients of loss function w.r.t the input image in [38]. Therefore, feature-level scores are no longer required. However, this strategy may not achieve satisfactory results for general methods in practice [31]. Finally, the work [40] proposed a refinement network to refine the initial anomaly map, utilizing the input image as additional channels. However, this method requires more samples to train the refinement network. In addition, the mixture of normal and anomalous pixels described in Issue 2 remains unsolved in the above methods [17, 38, 40].
@@ III-D Optimization Algorithm
According to our discussions in Section III-B and Eqs. (8), (10), and (11), it is possible to use the gradient-based algorithm to solve the problem (12). However, due to the complicated structure of (12) built on deep neural networks, the solutions of some pixels obtained by popular algorithms may be poor. The gradients will diminish at these pixels and thus early stop. To deal with this challenge, we propose a novel local gradient-sharing mechanism, as illustrated in Fig. 3.
@@ IV-C Ablation Study
Type F2PAD PatchCore F2PAD CFLOW-AD F2PAD PaDiM
@@ V Conclusion
Unsupervised image anomaly detection is critical for various manufacturing applications, where precise pixel-level anomaly segmentation is required for downstream tasks. However, existing feature-based methods show inaccurate anomaly boundaries due to the decreased resolution of the feature map and the mixture of adjacent normal and anomalous pixels. To address the above issues, we proposed a novel optimization framework, namely F2PAD, to enhance various feature-based methods, including the popular PatchCore, CFLOW-AD, and PaDiM. Our extensive case studies demonstrated that our F2PAD improved remarkably over three backbone methods in normal and few-shot settings. Moreover, the necessity of each component in F2PAD was demonstrated by our ablation studies. Our F2PAD is suitable to handle realistic and complex manufacturing scenarios, such as small-batch manufacturing, and has the potential to become a general framework to enhance feature-based anomaly detection methods.

### v2

@@ Abstract
Image-based inspection systems have been widely deployed in manufacturing production lines. Due to the scarcity of defective samples, unsupervised anomaly detection that only leverages normal samples during training to detect various defects is popular. Existing feature-based methods, utilizing deep features from pretrained neural networks, show their impressive performance in anomaly localization and the low demand for the sample size for training. However, the detected anomalous regions of these methods always exhibit inaccurate boundaries, which impedes the downstream tasks. This deficiency is caused: (i) The decreased resolution of high-level features compared with the original image, and (ii) The mixture of adjacent normal and anomalous pixels during feature extraction. To address them, we propose a novel unified optimization framework (F2PAD) that leverages the Feature-level information to guide the optimization process for Pixel-level Anomaly Detection in the inference stage. The proposed universal framework can enhance various feature-based methods with limited assumptions. Case studies are provided to demonstrate the effectiveness of our strategy, particularly when applied to three popular backbone methods: PaDiM, CFLOW-AD, and PatchCore. Our source code and data will be available upon acceptance.
@@
Image anomaly detection, unsupervised learning, anomaly segmentation, feature extraction, neural networks.
@@ I Introduction
Automated visual inspection of products by deploying cost-efficient image sensors has become a critical concern in industrial applications in recent years, ranging from microscopic fractions and crack detection on electrical commutators [1] to fabric defect detection on textiles [2]. In realistic production lines, defective products are always scarce. Moreover, all possible types of defects that occurred during production are impossible to know beforehand [3]. In this scenario, only utilizing normal (non-defective) samples to train modern machine learning models for defect detection becomes essential. Therefore, unsupervised anomaly detection has become a mainstream setting in practical applications.
Besides determining whether a sample is anomalous or not, it is also essential to precisely localize and segment the defective regions due to the following reasons: (i) It enables the characterization of surface defects, such as shape, size, and orientation, which contributes to the evaluation of product quality. (ii) It benefits the subsequent defect classification and recognition, as the models could focus more on the discriminative regions. An example is provided in Fig. 1, where the target defective region is masked as white in the ground truth (GT).
@@ II-A Image Reconstruction-Based Methods
Reconstruction-based image anomaly detection methods aim to reconstruct the normal samples accurately during training, assuming imperfect reconstruction of defective regions for an anomalous image during inference. The pixel differences between the input and reconstructed images can be directly transformed into anomaly scores and further thresholded as a binary anomaly mask. Therefore, the key research goal is to pursue a more accurate reconstruction. Autoencoder (AE) is the most prevalent architecture for image reconstruction [20, 21]. In AEs, a common choice of the backbone network is CNN. Vision transformers were introduced in [22, 23] to capture more global structural information. Besides the AEs, GANs [7, 24] and diffusion models [25, 26] are also popular choices.
Instead of focusing on the architecture of the reconstruction network, there are various strategies to improve the reconstruction accuracy. For example, the structure similarity index measure (SSIM) was introduced in [5] to incorporate the perceptual difference of anomaly. Additionally, to address the deterioration of image reconstruction caused by anomalies, [27, 28] enforced the reconstructed image on the learned manifold. Alternatively, [29] considered the correspondence between structure and texture.
@@ II-B Feature-Based Methods
PaDiM [9] adopted the multivariate Gaussian distributions to model the normal features. To allow more complicated probabilistic distributions, DifferNet [11] adopted a neural network, i.e., normalizing flow, to estimate the density of features. Furthermore, CFLOW-AD [12] improved the DifferNet by modeling the conditional probabilities of features on their locations. Similarly, a 2D normalizing flow was proposed in [30] to retain the spatial relationship between features, while the multi-scale features were considered in [31]. The memory bank-based method that collects all normal features was first proposed in [32], while the average distance of a test feature to its k-nearest neighbors was regarded as the anomaly score. To improve the inference speed, the coreset technique was adopted in PatchCore[16] by greedily selecting the farthest features to construct a reduced memory bank. Similarly, [33] proposed a gradient preference strategy to determine the most significant features for anomaly detection. In addition, [34] encouraged the normal features within compact spheres. Teacher-student networks were also popular in literature [35], which effectively transfer normal features from the teacher network to the student model but fail for anomalous features.
One issue of the above methods is the decreased resolution of the feature map, as stated in Issue 1. To tackle it, the work [13] learned dense features in the image size by using the multipooling and unwarping strategy in [36]. However, it is only applicable to local patch descriptions by CNNs. Furthermore, the anomaly scores were defined as the gradients of loss function w.r.t the input image in [35]. Therefore, feature-level scores are no longer required. However, this strategy may not achieve satisfactory results for general methods in practice [27]. Finally, the works [37] and [38] adopted additional segmentation networks to refine the initial anomaly map. However, these methods require more samples to train the segmentation networks. In addition, the mixture of normal and anomalous pixels described in Issue 2 remains unsolved in the above methods [13, 35, 37].
@@ III-D Optimization Algorithm
According to our discussions in Section III-B and Eqs. (8), (10), and (11), it is possible to use the gradient-based algorithm to solve the problem (12). However, due to the complicated structure of (12) built on deep neural networks, the solutions of some pixels obtained by popular algorithms may be poor. The gradients will diminish at these pixels and thus early stop. To address this challenge, we propose a novel local gradient-sharing mechanism, as illustrated in Fig. 3.
@@ IV-C Ablation Study
Ablation F2PAD PatchCore F2PAD CFLOW-AD F2PAD PaDiM
@@ IV-C3 Performance of Optimization Algorithm
IV-D Discussion
TABLE IV: The average number of iterations and running time per sample during optimization in the normal setting.
Item F2PAD PatchCore F2PAD CFLOW-AD F2PAD PaDiM
# Iteration 202.3 204.0 398.1
Time (Seconds) 17.8 8.7 40.8
IV-D1 Time Analysis
We summarize the running time of our F2PAD to demonstrate its practicability. Specifically, we implemented our F2PAD on Python 3.9 with CPU Intel i7 10700, GPU GTX 3080 10GB, and RAM 64GB. The average number of iterations until convergence and total time of solving the problem (12) in the normal setting in Section IV-B1 are reported in Table IV. With a minimal average total time per sample of 8.7 seconds for F2PAD CFLOW-AD and a maximum of 40.8 seconds for F2PAD PaDiM, our F2PAD demonstrates good applicability to real-world quality control scenarios. Notably, the relatively low computational speed of F2PAD PaDiM is related to the large number of $\{\mathbf{\mu}_{i,j}\}$ and $\{\mathbf{\Lambda}_{i,j}\}$ required for each location $i,j$ in Eq. (7). One may consider to approximate $\{\mathbf{\mu}_{i,j}\}$ and $\{\mathbf{\Lambda}_{i,j}\}$ with fewer bases, which remains our future study.
IV-D2 Failure Analysis
Figure 11: Visualization of failure scenarios. (a) Cases in which the initial masks, obtained through the backbone method, are significantly incomplete. (b) Cases in which specific regions of anomalies appear ambiguous, and the associated areas in the images recovered by our F2PAD exhibit properties akin to those of normal areas.
Fig. 11 summarizes two types of failure cases in our experiment. The first case in Fig. 11(a) stems from incomplete masks generated by backbone methods. To initialize our algorithm (see Section III-E2), we concentrate on areas within dilated masks, which may occasionally fail to cover the GTs. The second case, illustrated in Fig. 11(b), occurs when our F2PAD fails to recover some anomalous pixels. These pixels are visually compatible with the recovered images, yielding appearances nearly indistinguishable from normal ones. Nevertheless, these failed cases are rare in our experiment.
@@ V Conclusion
Unsupervised image anomaly detection is critical for various manufacturing applications, where precise pixel-level anomaly segmentation is required for downstream tasks. However, existing feature-based methods show inaccurate anomaly boundaries due to the decreased resolution of the feature map and the mixture of adjacent normal and anomalous pixels. To address the above issues, we proposed a novel optimization framework, namely F2PAD, to enhance various feature-based methods, including the popular PatchCore, CFLOW-AD, and PaDiM. Our extensive case studies demonstrated that our F2PAD improved remarkably over three backbone methods in normal and few-shot settings. Moreover, the necessity of each component in F2PAD was demonstrated by our ablation studies. Our F2PAD is suitable for handling realistic and complex manufacturing scenarios, such as small-batch manufacturing. It has the potential to become a general framework to enhance feature-based anomaly detection methods.